if you purchase 10 call contracts with strike price of $100 and a premium of $15 per share, what would your net dollar return be if the price rose to $125?

a) $25,000

b) $15,000

c) $12,500

d) $10,000
Answer #1

Profit on Call Option = Max(0, Stock Price - Strike Price) - Premium Paid

Profit on 1 call option = (125 - 100 - 15)(100) = $1,000

Profit on 10 call options = 10(1,000)

Profit on 10 call options = $10,000